OpenStudy (anonymous):

Write an equation of a line that is perpendicular to the line y = 5x + 1 and contains the point (-10,2).

OpenStudy (anonymous):

@Daniellelovee @Littlebird

OpenStudy (littlebird):

perpendicular lines have slopes that are negative reciprocals of eachother

OpenStudy (anonymous):

So y = - 1/5x + 1

OpenStudy (anonymous):

Do I have to change the y intercept like last time?

OpenStudy (littlebird):

Yes, because they want you to include a specific point

OpenStudy (anonymous):

52 = b? That feels wrong >.>

OpenStudy (anonymous):

I did 2 = 5(-10) + b 2 = -50 + b Add 50 to both sides 52 = b

OpenStudy (anonymous):

Put the 2 in for y and the -10 in for x

OpenStudy (littlebird):

Remember that you have to change the slope to -1/5

OpenStudy (anonymous):

y = -1/5x + 52

OpenStudy (littlebird):

No, before you find the y-intercept, change the slope

OpenStudy (anonymous):

OOOOH , my bad

OpenStudy (anonymous):

So it would: 2 = -1/5(-10) + b Then 2 = 2 + b Subtract 2 from both sides b = 0

OpenStudy (anonymous):

Then final answer would be: y = -1/5x + 0

OpenStudy (littlebird):

yes, but in your real answer, you shouldn't have to write out the + 0

OpenStudy (anonymous):

Yeah, I thought so ;P